### v1

@@ Abstract
Context:Two protoplanets have recently been discovered within the PDS 70 protoplanetary disk. JWST/NIRCam offers a unique opportunity to characterize them and their birth environment at wavelengths difficult to access from the ground.
Aims:We aim to image the circumstellar environment of PDS 70 at 1.87 m and 4.83 m, assess the presence of Pa- emission due to accretion onto the protoplanets, and probe any IR excess indicative of heated circumplanetary material.
Methods:We obtain non-coronagraphic JWST/NIRCam images of PDS 70 within the MINDS (MIRI mid-INfrared Disk Survey) program. We leverage the Vortex Image Processing (VIP) package for data reduction, and develop dedicated routines for optimal stellar PSF subtraction, unbiased imaging of the disk, and protoplanet flux measurement in this type of dataset. A radiative transfer model of the disk is used to disentangle the contributions from the disk and the protoplanets.
Results:We re-detect both protoplanets and identify extended emission after subtracting a disk model, including a large-scale spiral-like feature. We interpret its signal in the direct vicinity of planet as tracing the accretion stream feeding its circumplanetary disk, while the outer part of the feature may rather reflect asymmetric illumination of the outer disk. We also report a bright signal consistent with a previously proposed protoplanet candidate enshrouded in dust, near the 1:2:4 mean-motion resonance with planets and . The 1.87 m flux of planet is consistent with atmospheric model predictions, but not that of planet . We discuss potential origins for this discrepancy, including significant Pa- line emission. The 4.83 m fluxes of planets and suggest enshrouding dust or heated CO emission from their circumplanetary environment.
Conclusions:The use of image-processing methods optimized for extended disk signals on high-sensitivity and high-stability JWST images can uniquely identify signatures of planet–disk interactions and enable accurate photometry of protoplanets at wavelengths difficultly probed from the ground. Our results point towards the prospect of identifying and characterizing more protoplanets in other JWST datasets.
MINDS: JWST/NIRCam imaging of the protoplanetary disk PDS 70
@@
The direct detection and characterization of protoplanets is an emerging research field enabled by the latest generations of high-contrast and high-angular resolution instruments available on the largest ground-based facilities (e.g., Keppler et al. 2018; Hammond et al. 2023). While the application of post-processing algorithms designed for point-source detection to bright protoplanetary disks has led to a number of unconfirmed protoplanet candidates, the recent development of both iterative and inverse-problem image-processing approaches (IIPAs) is opening the way to unbiased high-contrast IR imaging of the birth environment of planets (e.g., Pairet et al. 2021; Flasseur et al. 2021; Juillard et al. 2023). JWST now offers the opportunity to use these optimized tools to deepen the search and characterization of protoplanets at wavelengths inaccessible or difficult to observe from the ground. This Letter demonstrates the combined potential of JWST and IIPAs, using as a testbed the PDS 70 system.
@@
PDS 70 is a young ( Myr) K7IV star located at a distance of pc (Müller et al. 2018; Gaia Collaboration et al. 2021). It is surrounded by a protoplanetary disk composed of a water-rich inner disk (Dong et al. 2012; Perotti et al. 2023), separated from the outer disk by an annular gap extending up to a radius of 54 au (Long et al. 2018; Keppler et al. 2019). Two nascent planets were imaged and confirmed independently in this large gap, at multiple NIR and sub-mm wavelengths, as well as in the H line filter (e.g., Keppler et al. 2018; Müller et al. 2018; Christiaens et al. 2019b; Haffert et al. 2019; Isella et al. 2019; Benisty et al. 2021). A third protoplanet candidate was also reported on a potential 13.5 au orbit (Mesa et al. 2019, hereafter M19). As such, the system is an ideal laboratory to study planet-disk interactions and search for accretion signatures. Hydro-dynamical simulations suggest that the large gap is dynamically carved by both planets, with the (near) 2:1 mean-motion resonance observed by GRAVITY (Wang et al. 2021b) explained as the outcome of planet migration followed by resonance capture (Bae et al. 2019; Toci et al. 2020). Constraints on the distribution and depletion of dust and gas in the disk were recently inferred through radiative transfer and thermo-chemical forward modeling of the NIR and sub-mm observations (Portilla-Revelo et al. 2022, 2023). Both sub-mm continuum and NIR observations have also identified an arm-like structure hypothesized to trace either an asymmetric outer ring or a gap-induced vortex (Isella et al. 2019; Juillard et al. 2022).
@@
Spectral characterization of protoplanets and favors dust-enshrouded atmospheres combined with a potential IR excess (Müller et al. 2018; Christiaens et al. 2019a; Wang et al. 2020; Stolker et al. 2020; Wang et al. 2021b). Yet, cross-correlation of either these best-fit models or molecular templates with medium-resolution spectra did not detect the planets (Cugno et al. 2021). This suggests that either significantly more extinction affects both protoplanets than SED fitting implies, or that self-consistent atmospheric models used to characterize adolescent sub-stellar objects do not describe embedded protoplanets. Either hybrid models with surface accretion shock contributions (e.g., Aoyama et al. 2020, 2021), or circumplanetary disk models, may more accurately account for the measured SED (e.g., Chen & Szulágyi 2022). In this context, observations at additional wavelengths are key to disentangle the different models.
@@
Notes: Radial separation. Position angle measured east of north. Contrast ratio with respect to the star and unresolved inner disk.
@@ 2 Observations and image processing
We observed PDS 70 with JWST/NIRCam (Rieke et al. 2005) as part of the MIRI Guaranteed Time Observations on protoplanetary disks MINDS survey (PI: Th. Henning, PID: 1282) on 8 March 2023. Images were acquired simultaneously in the F187N ( m, m) and F480M ( m, m) filters without coronagraph, with pixel scales of 31 and 63 mas/pixel, respectively. We used the smallest subarray (SUB64P), with an effective integration time of 0.35112 s, 7 groups per integration, 142 integrations per exposure, and 5 dithered exposures, each at two roll angles separated by 5.0°. The total effective integration time was thus 8.3 min.
We calibrated the data with stages 1 and 2 (Detector1 and Image2) of the JWST pipeline (v1.10.2; Bushouse et al. 2023), using Calibration Reference Data System context jwst_1166.pmap. The calibration yielded two sets of five dithered images, each set corresponding to a different roll angle. We then relied on routines from the Vortex Image Processing (VIP; Gomez Gonzalez et al. 2017; Christiaens et al. 2023) package for image pre-processing (details in Appendix A.1), stellar PSF subtraction, and forward modeling of circumstellar signals. As the star did not saturate the detector, we performed aperture photometry directly on the pre-processed images. We set the largest aperture radius fitting within the SUB64P field of view. Table 1 reports the integrated photometry of the system. We conservatively consider 1% absolute flux calibration uncertainties.
We investigated different approaches for PSF modeling and subtraction (details in Appendix A.2). The best-quality images of the system were obtained with our implementation of Iterative Principal Component Analysis (IPCA; e.g., Pairet et al. 2021; Stapper & Ginski 2022) leveraging roll angle diversity. Its principle relies on estimating circumstellar signals in the processed image obtained at each iteration, and removing them from the images used to create a PSF model (i.e., for each image, the ones obtained at the other roll angle) in the next iteration. We applied IPCA to a test dataset to illustrate its reliability at recovering point-like and extended circumstellar signals (Appendix A.3).
Before extracting the photometry of the protoplanets, one needs to remove the expected contribution of the disk. This is particularly relevant for planet c, located near the bright edge of the outer disk. We rely on the latest radiative transfer models of the disk detailed in Portilla-Revelo et al. (2022, 2023), and followed a procedure which we refer to as the Negative Fake Disk technique (NEGFD; details in Appendix B). After removing the disk contribution, we used the Negative Fake Companion (NEGFC; e.g., Lagrange et al. 2010) technique to extract the exact astrometry and photometry of the protoplanets in both datasets (details in Appendix C). Our final retrieved astrometry and photometry for the protoplanets are presented in Table 1.
@@ 3.1 Spiral accretion stream or variable illumination?
Figure 1a and d show our F187N and F480M images obtained with IPCA, respectively. IPCA recovered faint circumstellar signals originally hidden in the wings of the stellar PSF, while iteratively correcting for geometric biases affecting extended signals in roll-subtracted images (Fig. 4a and e; Juillard et al. 2022). The IPCA images reveal inner and outer disk signals, signals at the expected location of the protoplanets, and a shift of the maximum intensity of the outer disk to the north of the semi-minor axis on the near side of the disk (located at a position angle east of north PA). In the absence of disk asymmetry or protoplanets, the maximum in total intensity is otherwise expected to be located along PA, based on the scattering phase function of sub-m size dust grains (e.g., Milli et al. 2017). To highlight the protoplanets and any disk asymmetry, we subtracted the disk model found with NEGFD from the IPCA images (Fig. 1c and f). Doing so reveals residual extended signals in addition to protoplanets and c, the predicted locations of which are indicated for the epoch of the observations. These predictions are based on the orbital fits presented in Wang et al. (2021b), and available through the platform whereistheplanet (Wang et al. 2021a). We do not detect any significant counterpart for the proposed sub-mm continuum signal at the L5 Lagrangian point associated with planet (Balsalobre-Ruza et al. 2023). We note however an outstanding extended spiral-like signal connected with the position of planet , indicated with a thick arrow in Fig. 1b and e.
@@
Dynamical interaction between a protoplanet and the disk in which it is embedded has long been known to cause spiral density waves (Goldreich & Tremaine 1979; Ogilvie & Lubow 2002) and angular momentum transport (Lin & Papaloizou 1979; Rafikov 2002). In the vicinity of the planet, the latter is expected to lead to a spiral-shaped accretion stream (Lubow et al. 1999; Ayliffe & Bate 2009). The accretion stream associated to PDS 70 was predicted in a dedicated 3D hydro-dynamical simulation in Toci et al. (2020). Figure 10 compares it to the spiral-like signal identified in our observations. Although the agreement is remarkable, we note that the accretion stream is mostly delimiting the edge of the cavity. An asymmetric illumination of the outer disk edge, not captured by our radiative transfer disk model, could equally lead to excess signal in our images after disk model subtraction. Multi-epoch polarised intensity images of the system show varying illumination and shadowing effects on the northwest and southeast parts of the outer disk (e.g., Fig. A2 in Juillard et al. 2022), suggesting this is likely the dominant cause for the observed excess. This interpretation is consistent with the significant mid-IR SED variability measured for the system, also suggesting variable shadowing effects from the inner parts onto the outer parts of the disk (Perotti et al. 2023).
@@
While illumination effects are the most likely culprit for excess signals at the edge of the outer disk (upper arc in Fig. 1c), such effects alone cannot account for the part of the spiral-like signal located inside the cavity, given the level of dust depletion therein (e.g., Dong et al. 2012; Keppler et al. 2018). In the direct vicinity of planet (lower arc), a genuine dust density enhancement appears thus necessary. Given the clear visual connection to the location of protoplanet , a spiral accretion stream feeding the circumplanetary disk of planet c, detected in sub-mm continuum observations (Isella et al. 2019; Benisty et al. 2021; Casassus & Cárcamo 2022), is the most straightforward explanation for this signal. Excess signal has tentatively been observed there in IIPA-processed VLT/SPHERE images (Flasseur et al. 2021; Juillard et al. 2022), although the non-removal of an outer disk model complicates its unambiguous identification in those images. Moreover, the location of the accretion stream is coincident with a gap-crossing spur found in ALMA CO observations of the disk (Keppler et al. 2019), further strengthening the accretion stream interpretation. This result suggests that some of the observed spiral features in less inclined disks than PDS 70 could also be associated with embedded protoplanets (e.g., Dong et al. 2018; Ren et al. 2024). A similar spiral-shaped signal was recently identified as connected to HD 169142 (Hammond et al. 2023). Likewise, a gap-crossing filament coincident with a twist in one of the spirals of HD 135344 B may also be caused by an embedded protoplanet (Casassus et al. 2021).
@@
An arm-like feature was identified and characterized in the outer disk of PDS 70 using multi-epoch SPHERE images of the system obtained with IIPAs (Pairet et al. 2021; Juillard et al. 2022). The trace of this arm in a 2021 SPHERE dataset is compared to our IPCA image in Fig. 10c. It is unclear whether it is associated with the spiral accretion stream, or whether it traces a separate feature, such as a vortex or an asymmetric second ring (Juillard et al. 2022). The inner part of the arm, at the edge of the cavity, appears consistent with the part of the spiral-like signal seen in the F187N image which may trace varying illumination of the outer disk edge. The outer part of the arm appears consistent with the arm-like feature identified in previous SPHERE images. This feature is the only signal detected at a SNR in our F480M images besides the protoplanets (Fig. 11), and was also observed in Keck/NIRC2 images of the system after a similar NEGFD procedure as ours (Fig. 1 in Wang et al. 2020). Since an alternative origin for the arm-like signal is the presence of a yet undetected planet in the outer disk exciting an inner spiral wake, we investigated the sensitivity of our data to additional planets (Appendix E). Our F480M contrast and corresponding mass sensitivity curves (Fig. 12) constrain any planets in the outer disk to have a mass below 1–2 , neglecting extinction.

### v2

@@ Abstract
Context. Two protoplanets have recently been discovered within the PDS 70 protoplanetary disk. JWST/NIRCam offers a unique opportunity to characterize them and their birth environment at wavelengths that are difficult to access from the ground.
Aims. We image the circumstellar environment of PDS 70 at 1.87 m and 4.83 m, assess the presence of Pa- emission due to accretion onto the protoplanets, and probe any IR excess indicative of heated circumplanetary material.
Methods. We obtained noncoronagraphic JWST/NIRCam images of PDS 70 within the MIRI mid-INfrared Disk Survey (MINDS) program. We leveraged the Vortex Image Processing (VIP) package for data reduction, and we developed dedicated routines for optimal stellar point spread function subtraction, unbiased imaging of the disk, and protoplanet flux measurement in this type of dataset. A radiative transfer model of the disk was used to separate the contributions from the disk and the protoplanets.
Results. We redetect both protoplanets and identify extended emission after subtracting a disk model, including a large-scale spiral-like feature. We interpret its signal in the direct vicinity of planet as tracing the accretion stream that feeds its circumplanetary disk, while the outer part of the feature may rather reflect asymmetric illumination of the outer disk. We also report a bright signal that is consistent with a previously proposed protoplanet candidate enshrouded in dust near the 1:2:4 mean-motion resonance with planets and . The 1.87 m flux of planet is consistent with atmospheric model predictions, but the flux of planet is not. We discuss potential origins for this discrepancy, including significant Pa- line emission. The 4.83 m fluxes of planets and suggest enshrouding dust or heated CO emission from their circumplanetary environment.
Conclusions. The use of image-processing methods that are optimized for extended disk signals on high-sensitivity and high-stability from JWST can uniquely identify signatures of planet–disk interactions and enable accurate photometry of protoplanets at wavelengths that are difficult to probe from the ground. Our results indicate that more protoplanets can be identified and characterized in other JWST datasets.
@@
The direct detection and characterization of protoplanets is an emerging research field that is enabled by the latest generations of instruments achieving high contrast and high angular resolution that are available on the largest ground-based facilities (e.g., Keppler et al. 2018; Hammond et al. 2023). While the application of post-processing algorithms designed for point-source detection to bright protoplanetary disks has led to a number of unconfirmed protoplanet candidates, the recent development of both iterative and inverse-problem image-processing approaches (IIPAs) is opening the way to unbiased high-contrast IR imaging of the birth environment of planets (e.g., Pairet et al. 2021; Flasseur et al. 2021; Juillard et al. 2023). JWST now offers the opportunity to use these optimized tools to deepen the search and characterization of protoplanets at wavelengths that are inaccessible or are difficult to observe from the ground. This Letter demonstrates the combined potential of JWST and IIPAs and uses the PDS 70 system as a testbed.
@@
PDS 70 is a young ( Myr) K7IV star located at a distance of pc (Müller et al. 2018; Gaia Collaboration et al. 2021). It is surrounded by a protoplanetary disk composed of a water-rich inner disk (Dong et al. 2012; Perotti et al. 2023) that is separated from the outer disk by an annular gap extending up to a radius of 54 au (Long et al. 2018; Keppler et al. 2019). Two nascent planets were imaged and confirmed independently in this large gap, at multiple near-IR (NIR) and submillimeter (submm) wavelengths, as well as in the H line filter (e.g., Keppler et al. 2018; Müller et al. 2018; Christiaens et al. 2019b; Haffert et al. 2019; Isella et al. 2019; Benisty et al. 2021). A third protoplanet candidate was also reported on a potential 13.5 au orbit (Mesa et al. 2019, hereafter M19). The system is therefore an ideal laboratory in which to study planet-disk interactions and search for accretion signatures. Hydrodynamical simulations suggest that the large gap is dynamically carved by both planets, and the (near) 2:1 mean-motion resonance observed by GRAVITY (Wang et al. 2021b) is explained as the outcome of planet migration followed by resonance capture (Bae et al. 2019; Toci et al. 2020). Constraints on the distribution and depletion of dust and gas in the disk were recently inferred through radiative transfer and thermo-chemical forward modeling of the NIR and submm observations (Portilla-Revelo et al. 2022, 2023). Submm continuum and NIR observations also identified an arm-like structure that is hypothesized to either trace an asymmetric outer ring or a gap-induced vortex (Isella et al. 2019; Juillard et al. 2022).
@@
The spectral characterization of protoplanets and favors dust-enshrouded atmospheres combined with potential IR excess (Müller et al. 2018; Christiaens et al. 2019a; Wang et al. 2020; Stolker et al. 2020; Wang et al. 2021b). The cross-correlation of either these best-fit models or molecular templates with medium-resolution spectra did not detect the planets, however (Cugno et al. 2021). This suggests that either significantly more extinction affects the two protoplanets than fitting of the spectral energy distribution (SED) implies or that the self-consistent atmospheric models that have been used to characterize adolescent substellar objects do not describe embedded protoplanets. Hybrid models with contributions from surface accretion shocks (e.g., Aoyama et al. 2020, 2021), or circumplanetary disk models may more accurately account for the measured SED (e.g., Chen & Szulágyi 2022). In this context, observations at additional wavelengths are key to distinguish the different models.
@@
Notes: (a)Radial separation. (b)Position angle measured east of north. (c)Contrast ratio with respect to the star and unresolved inner disk.
@@ 2 Observations and image processing
We observed PDS 70 with JWST/NIRCam (Rieke et al. 2005) as part of the MIRI guaranteed time observations on protoplanetary disks in the MINDS survey (PI: Th. Henning, PID: 1282) on 8 March 2023. The images were acquired simultaneously in the F187N ( m, m) and F480M ( m, m) filters without the coronagraph, with pixel scales of 31 and 63 mas/pixel, respectively. We used the smallest subarray (SUB64P), with an effective integration time of 0.35112 s, seven groups per integration, 142 integrations per exposure, and five dithered exposures at two roll angles each, separated by 5.0°. The total effective integration time was thus 8.3 min.
We calibrated the data with stages 1 and 2 (Detector1 and Image2) of the JWST pipeline (v1.10.2; Bushouse et al. 2023), using the calibration reference data system context jwst_1166.pmap. The calibration yielded two sets of five dithered images, each set corresponding to a different roll angle. We then relied on routines from the Vortex Image Processing (VIP; Gomez Gonzalez et al. 2017; Christiaens et al. 2023) package for image preprocessing (details in Appendix A.1), stellar PSF subtraction, and forward modeling of circumstellar signals. As the star did not saturate the detector, we directly performed aperture photometry on the preprocessed images. We set the largest aperture radius that fit within the SUB64P field of view. Table 1 reports the integrated photometry of the system. We conservatively considered 1% absolute flux calibration uncertainties.
We investigated different approaches for PSF modeling and subtraction (details in Appendix A.2). The best-quality images of the system were obtained with our implementation of an iterative principal component analysis (IPCA; e.g., Pairet et al. 2021; Stapper & Ginski 2022) that leverages roll angle diversity. Its principle relies on estimating the circumstellar signals in the processed image that are obtained at each iteration and removing them from the images that are used to create a PSF model (i.e., for each image, those obtained at the other roll angle) in the next iteration. We applied IPCA to a test dataset to illustrate that it reliably recovers point-like and extended circumstellar signals (Appendix A.3).
Before the photometry of the protoplanets is extracted, the expected contribution of the disk needs to be removed. This is particularly relevant for planet c, which is located near the bright edge of the outer disk. We relied on the latest radiative transfer models of the disk detailed in Portilla-Revelo et al. (2022, 2023), and followed a procedure that we refer to as the negative fake disk technique (NEGFD; details in Appendix B). After removing the disk contribution, we used the negative fake companion (NEGFC; e.g., Lagrange et al. 2010) technique to extract the exact astrometry and photometry of the protoplanets in both datasets (details in Appendix C). The final astrometry and photometry we retrieved for the protoplanets are presented in Table 1.
@@ 3.1 Spiral accretion stream or variable illumination?
Panels a and e in Figure 1 show the F187N and F480M images we obtained with IPCA, respectively. IPCA recovered faint circumstellar signals that were originally hidden in the wings of the stellar PSF while it iteratively corrected for geometric biases that affect extended signals in roll-subtracted images (Fig. 4a and e; Juillard et al. 2022). The IPCA images reveal signals of the inner and outer disk, signals at the expected location of the protoplanets, and a shift in the maximum intensity of the outer disk north of the semiminor axis on the near side of the disk (located at a position angle east of north PAβ). In the absence of disk asymmetry or protoplanets, the maximum in total intensity is otherwise expected to be located along PAβ, based on the scattering phase function of sub-m size dust grains (e.g., Milli et al. 2017). To highlight the protoplanets and any disk asymmetry, we subtracted the disk model found with NEGFD from the IPCA images (Fig. 1c and f). This revealed residual extended signals in addition to protoplanets and c, the predicted locations of which are indicated for the epoch of the observations. These predictions are based on the orbital fits presented in Wang et al. (2021b), and they are available through the platform whereistheplanet (Wang et al. 2021a). We do not detect any significant counterpart for the proposed submm continuum signal at the L5 Lagrangian point associated with planet (Balsalobre-Ruza et al. 2023). We note an outstanding extended spiral-like signal connected with the position of planet , however. It is indicated with a thick arrow in Fig. 1b and e.
@@
The dynamical interaction between a protoplanet and the disk in which it is embedded has long been known to cause spiral density waves (Goldreich & Tremaine 1979; Ogilvie & Lubow 2002) and angular momentum transport (Lin & Papaloizou 1979; Rafikov 2002). In the vicinity of the planet, the latter is expected to lead to a spiral-shaped accretion stream (Lubow et al. 1999; Ayliffe & Bate 2009). The accretion stream associated with PDS 70 was predicted in a dedicated 3D hydrodynamical simulation in Toci et al. (2020). Figure 10 compares it with the spiral-like signal we identified in our observations. Although the agreement is remarkable, we note that the accretion stream mostly delimits the edge of the cavity. An asymmetric illumination of the outer-disk edge that is not captured by our radiative transfer disk model could equally lead to an excess signal in our images after disk model subtraction. Multi-epoch polarized intensity images of the system show varying illumination and shadowing effects in the northwest and southeast parts of the outer disk (e.g., Fig. A2 in Juillard et al. 2022), suggesting that this is likely the dominant cause for the observed excess. This interpretation is consistent with the significant mid-IR (MIR) SED variability measured for the system, which also suggests variable shadowing effects from the inner parts onto the outer parts of the disk (Perotti et al. 2023).
@@
While illumination effects most likely cause excess signals at the edge of the outer disk (upper arc in Fig. 1c), these effects alone cannot account for the part of the spiral-like signal that is located inside the cavity based on the level of dust depletion therein (e.g., Dong et al. 2012; Keppler et al. 2018). In the direct vicinity of planet (lower arc), a genuine dust density enhancement therefore appears to be required. Based on the clear visual connection to the location of protoplanet , a spiral accretion stream feeding the circumplanetary disk of planet c that is detected in submm continuum observations (Isella et al. 2019; Benisty et al. 2021; Casassus & Cárcamo 2022) is the most straightforward explanation for this signal. Excess signal has tentatively been observed there in IIPA-processed VLT/SPHERE images (Flasseur et al. 2021; Juillard et al. 2022), although the nonremoval of an outer-disk model complicates an unambiguous identification of this signal in these images. Moreover, the location of the accretion stream is coincident with a gap-crossing spur found in ALMA CO observations of the disk (Keppler et al. 2019). This further strengthens the interpretation that this is an accretion stream. This result suggests that some of the observed spiral features in less strongly inclined disks than PDS 70 could also be associated with embedded protoplanets (e.g., Dong et al. 2018; Ren et al. 2024). A similar spiral-shaped signal has recently been identified as connected to HD 169142 (Hammond et al. 2023). Likewise, a gap-crossing filament coincident with a twist in one of the spirals of HD 135344 B may also be caused by an embedded protoplanet (Casassus et al. 2021).
@@
An arm-like feature was identified and characterized in the outer disk of PDS 70 using multi-epoch SPHERE images of the system obtained with IIPAs (Pairet et al. 2021; Juillard et al. 2022). The trace of this arm in a 2021 SPHERE dataset is compared with our IPCA image in Fig. 10c. It is unclear whether it is associated with the spiral accretion stream or traces a separate feature, such as a vortex or an asymmetric second ring (Juillard et al. 2022). The inner part of the arm at the edge of the cavity appears to be consistent with the part of the spiral-like signal seen in the F187N image, which may trace the varying illumination of the outer-disk edge. The outer part of the arm appears to be consistent with the arm-like feature identified in previous SPHERE images. This feature is the only signal that is detected at a signal-to-noise ratio (S/N) in our F480M images in addition to the protoplanets (Fig. 11) and was also observed in Keck/NIRC2 images of the system after a similar NEGFD procedure as we used (Fig. 1 in Wang et al. 2020). An alternative origin for the arm-like signal is the presence of an as yet undetected planet in the outer disk that excites an inner spiral wake. We therefore investigated whether our data are sensitive to additional planets (Appendix E). Our F480M contrast and corresponding mass-sensitivity curves (Fig. 12) constrain any planets in the outer disk to have a mass below 1–2 , neglecting extinction.